Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 02 and week 03

Figure 1: Section 2.3, problem 43: mistake 1 Figure 2: Section 2.3, problem 43: mistake 2

2 Figure 3: Section 2.3, problem 43: mistake 3

3 Figure 4: Section 2.3, problem 43: mistake 4

4 Figure 5: Answer to homework week 03, problem 2

5